HELP HELP HELP!!! The formula S = C(1 + r)^t models inflation, where C = the value today, r = the annual inflation rate, and S = the inflated value t years from now. Use this formula to solve and round answer to nearest dollar amount.
If the inflation rate is 3%, how much will a house now worth $110,000 be worth in 5 years?

Answer by dolly(163)

S = C(1+r)^t
Here C = 110000, r = 0.03 , t = 5
So S = 110000(1+0.03)^5
= 110000(1.03)^5
= 110000(1.1592)
= 127512
So the house is worth $127512 in 5 years